OpenStudy (anonymous):

Anybody know how to do this?Medal will be given! At an all-you-can-eat barbecue fundraiser that you are sponsoring, adults pay 6$ for a dinner and children pay 4$ for a dinner. 212 people attend and raise 1128$. What is the total number of adults and the total number of children attending. (using systems of equations. a. what is a system of equations that you can solve this problem? b.what method would you use and why?

OpenStudy (maddieprough):

Try dividing 212 from $1,128 and then add 6 and 4 to that number evenly distibuted :)

OpenStudy (mathmale):

Megan: I responded earlier and gave you hints on how to get started. Were you not able to use that information? Why re-post your math problem?

OpenStudy (anonymous):

@maddieprough I have to use systems of equations to solve

OpenStudy (anonymous):

@mathmale I didn't get what you were saying so I just reposted it and btw i go by meg.

OpenStudy (anonymous):

Hey!!! We have the same name! Anyways, let's get started on your question! We need to make a system of equalities!

OpenStudy (mathmale):

If "a" adults pay $65 each for dinner, what will the total revenue be (expressed algebraically, not with numbers)? If "c" children pay $45 each, what will be the total revenue?

OpenStudy (anonymous):

system of equations**

OpenStudy (anonymous):

yes we do, but i go by meg.

OpenStudy (mathmale):

If there are "a" adults and "c" children, how would you express the total number of people at the dinner, algebraically?

OpenStudy (anonymous):

Like I said I thought it was x + y = 212 and 6x and 4y = 1128

OpenStudy (anonymous):

Ok, meg! We need to make 2 equations from the information you were given! The two equations you made are perfect!

OpenStudy (anonymous):

There right?

OpenStudy (anonymous):

why would you chnage it to that? @mathmale

OpenStudy (anonymous):

x + y = 212 6x + 4y = 1128 So, we just need to keep in mind that x = adults y = children

OpenStudy (anonymous):

I like solving by elimnation so... -4x -4y = -848 6x + 4y = 1128

OpenStudy (anonymous):

2x=280

OpenStudy (anonymous):

x = 140?

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

140+ y = 212 -140 -140 y = 72?

OpenStudy (anonymous):

And correct again, Meg! :D

OpenStudy (anonymous):

Thanks! just like reassurance on these problems!

OpenStudy (anonymous):

Great job! :)